OpenStudy (anonymous):

Express this amount in dozens. (Use decimal form.) 7.5 cupcakes

OpenStudy (anonymous):

dozens means sets of 12. For example 6 is half a dozen: 6/12 = 1/2 = 0.5 So 7.5 cupcakes in dozens is 7.5/12 = 15/24 = 5/8 dozen cupcakes = 0.625 dozen cupcakes

OpenStudy (anonymous):

7.5 / 12 = 0.625

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Made a type there. Should be (7.5) / (12) not (7/5) / (12). Final answer is .625

OpenStudy (anonymous):

1 dozen =12 then 7.5/12=0.63(converting to 2 decimals)